# AmazUtah_NLP at SemEval-2024 Task 9: A MultiChoice Question Answering System for Commonsense Defying Reasoning

### v2

@@ Abstract
The SemEval 2024 BRAINTEASER task represents a pioneering venture in Natural Language Processing (NLP) by focusing on lateral thinking, a dimension of cognitive reasoning that is often overlooked in traditional linguistic analyses. This challenge comprises of Sentence Puzzle and Word Puzzle subtasks and aims to test language models’ capacity for divergent thinking.
In this paper, we present our approach to the BRAINTEASER task. We employ a holistic strategy by leveraging cutting-edge pre-trained models in multiple choice architecture, and diversify the training data with Sentence and Word Puzzle datasets. To gain further improvement, we fine-tuned the model with synthetic humor/jokes dataset and the RiddleSense dataset which helped augmenting the model’s lateral thinking abilities. Empirical results show that our approach achieve 92.5% accuracy in Sentence Puzzle subtask and 80.2% accuracy in Word Puzzle subtask.
AmazUtah_NLP at SemEval-2024 Task 9: A MultiChoice Question Answering System for Commonsense Defying Reasoning
@@ 1 Introduction
The success of language models has inspired the Natural Language Processing community to attend to tasks that require implicit and complex reasoning. Human reasoning encompasses two types of reasoning: lateral and vertical thinking approaches. Lateral thinking demand out-of-the-box thinking. It is a form of creative reasoning that deviates from traditional, logical processes and has received little attention from NLP community. Vertical thinking on the other hand, relies on logical reasoning, and have been relatively popular in the past few years.
The BRAINTEASER dataset by (Jiang et al., 2023) stands as a crucial benchmark for evaluating question-answering systems. It particularly assesses these systems on their ability for lateral thinking — pushing them to transcend conventional commonsense reasoning towards more innovative and creative approaches to problem-solving. Additionally, as part of this effort to test language models’ lateral thinking capabilities, the SemEval 2024 BRAINTEASER task (Jiang et al., 2024), offers a focused challenge derived from the broader dataset, further probing the creative reasoning abilities of these models. This task is crucial because it addresses a gap in Natural Language Processing (NLP) where most tasks focus on linear, logical (vertical) thinking, neglecting the complex, divergent aspects of human cognition represented by lateral thinking. By encompassing two subtasks — Sentence Puzzle and Word Puzzle — BRAINTEASER aims to test a model’s ability to go beyond conventional commonsense associations, requiring an understanding of both standard meanings and the ability to reinterpret them in novel ways. This is vital for advancing the field of NLP, as it pushes the boundaries of what artificial intelligence can achieve in terms of mimicking the nuanced and creative aspects of human thought.
@@
Our system adopts a multifaceted strategy for this challenge, centering on the use of advanced pre-trained models like BERT (Devlin et al., 2019) and DeBERTaV3 (He et al., 2023) through HuggingFace’s (Wolf et al., 2020) AutoModelForMultipleChoice and AutoModelForSequenceClassification. This approach is enhanced by a diverse training regimen that mixes Sentence and Word Puzzle datasets, ensuring a broad exposure to different types of lateral thinking challenges. Additionally, the model is fine-tuned with a humor/jokes dataset generated by GPT-4 (OpenAI et al., 2024) and the RiddleSense (Lin et al., 2021) dataset, which introduces elements of creativity, unconventional thinking, and complex puzzle-solving. This comprehensive strategy aims to equip the model with enhanced lateral thinking abilities, crucial for tackling the creative and nuanced demands of the BRAINTEASER task in SemEval 2024.
In our participation in the BRAINTEASER task, we discovered that our system, particularly when finetuned with AutoModelForMultipleChoice, outperformed the baseline instruction-tuned systems mentioned in the original paper. This approach demonstrated a significant advantage in handling multiple-choice tasks. However, we faced challenges with AutoModelForSequenceClassification, suggesting an area for improvement. The incorporation of additional synthetic data and open-source dataset like RiddleSense positively influenced our performance. Quantitatively, our system achieved a commendable 6th place in the Sentence Puzzle and 10th in the Word Puzzle, indicating stronger proficiency in sentence-based challenges and room for growth in word-based puzzles.
@@
Our code and data will be available at https://github.com/soumyasmruti/semeval-2024-brainteaser after cleaning and de-anonimization.
@@ 3 Related Work
Researchers have explored various approaches to endow LLMs with commonsense reasoning abilities (Rae et al., 2021). One prominent approach is the use of knowledge graphs, which represent structured knowledge in the form of entities and their relationships (Ilievski et al., 2021). Authors in (Wang et al., 2021) proposed a method for incorporating commonsense knowledge from ConceptNet (Speer et al., 2018) into language models, leading to improved performance on commonsense reasoning tasks.
Another approach involves fine-tuning pre-trained LLMs on commonsense reasoning datasets. Authors in this paper (Huang et al., 2019) introduced the COSMOS QA dataset, which consists of multiple-choice questions that require commonsense reasoning. They showed that fine-tuning pre-trained LLMs on this dataset can significantly improve their commonsense reasoning capabilities.
Researchers have also investigated the use of prompting techniques to elicit commonsense reasoning from LLMs without explicit fine-tuning. (Zhou et al., 2022) proposed a method called "Conditional Prompt-Tuning" that enables LLMs to perform commonsense reasoning by conditioning on carefully designed prompts. In another work (Wei et al., 2022), chain-of-thought prompting showed how to unlock LLM’s reasoning ability via effective prompting techniques.
There hasn’t been extensive research on ‘lateral thinking‘ of LLMs. Very recenlty, OlaGPT (Xie et al., 2023) proposed a cognitive architecture framework in which they summarize various methods of human reasoning into Chain-of-Thought (CoT) templates, to maximize the LLMs’ reasoning effect.
Overall, while LLMs have shown flashes of non-linear, exploratory thinking on some benchmarks, lateral thinking as a holistic cognitive process remains an open challenge.
@@ 4.2 MultipleChoice QA with BERT and DeBERTa
4.2.1 Augmenting with RiddleSence and Humor Data
@@
The Ridlesense dataset consists of Riddles which are a form of puzzle where a question, often presented in a cryptic or metaphorical manner, challenges the reader to find a clever or unexpected answer.
To create the humor style QA, we prompted GPT 4. Crafting jokes content often requires a touch of ingenuity, an out-of-the-box approach, and a healthy dose of lateral thinking, and GPT-4 allowed us to explore unconventional and amusing angles to questions and answers. It’s like having a comedy writer who never runs out of fresh and unexpected punchlines. The details about how the dataset was generate is provided in Appendix  A.
@@ 6.1 Subtask A : Sentence Puzzle
Initially, we trained our models only on the provided sentence puzzle dataset but soon realized that combining both the sentence puzzle and word puzzle datasets yielded better validation scores. Consequently, we used the bert-base model with AutoModelForSequenceClassification, achieving an overall accuracy of 50.8%. Given that the dataset is in a multiple-choice format, we experimented with AutoModelForMultipleChoice using the same bert model. This change significantly improved performance, increasing accuracy by 10 points to 60%. Encouraged by this, we opted for the larger DeBERTaV3 model under the AutoModelForMultipleChoice configuration. This model, combined with the original dataset, significantly boosted performance, raising overall accuracy to 83.3%. After incorporating additional datasets containing humor-style questions and the RiddleSense dataset, our best accuracy score reached 92.5%. Our approach ranked 6th among the 31 teams that participated in the task and outperformed the baseline zero shot ChatGPT by almost 50 percentage points.
@@ 6.2 Subtask B : Word Puzzle
The word puzzle setup followed almost the same approach as sentence puzzle but during validation process we found the best model was the one which was trained with only original training data from word puzzle dataset and adding humor dataset. Adding RiddleSense data and sentence pzzle data didn’t improve the score of the word puzzle in validation process, therefore we didn’t submit that output. Our approach for this subtask didn’t perform that well when compared to other teams, we ranked 10th among the 23 teams that participated in this task, but outperformed the baseline zero shot ChatGPT by almost 40 percentage points.
@@ 7 Conclusion
In this work, we present our novel system designed for the SemEval 2024 BRAINTEASER task, which notably achieved 6th place in the Sentence Puzzle and 10th in the Word Puzzle categories. Our approach leverages advanced pre-trained models like BERT and DeBERTa, optimized through HuggingFace’s AutoModelForMultipleChoice and AutoModelForSequenceClassification. This strategy was further enhanced by incorporating a diverse training regimen, blending Sentence and Word Puzzle datasets with a unique humor/jokes dataset and the RiddleSense dataset. This mix has been instrumental in equipping our model with the lateral thinking capabilities essential for this task. While our system excelled in the Sentence Puzzle, reflecting a stronger grasp in sentence-based lateral thinking, the performance in the Word Puzzle highlighted areas for improvement, particularly in word-based lateral reasoning. The additional challenge posed by adversarial versions of puzzles, involving both Semantic and Context Reconstruction, underscores the complexity of this task. Our system’s performance underscores the efficacy of our training approach in enhancing lateral thinking in language models, a significant step forward in NLP. Future work will focus on refining our approach for word-based puzzles and further enhancing the model’s ability to navigate complex, creative reasoning paths, thereby advancing the field’s understanding of AI’s potential in mimicking nuanced aspects of human cognition.
@@ Appendix A Humor Dataset Details
We used the following prompts to generate Jokes or Humor style dataset. We experimented with multiple prompts and gather all the output in a json file and analyzed them manually.
@@
We initially prompted GPT 4 to generate 200 questions at once but that didn’t go well. The output contained duplicate questions after 15 / 16 unique ones. Basically, the model kept repeating itself. So we mostly used PROMPT 3 multiple times to generate high quality data. Before adding each we checked for duplicates again manually. Provided below are some of the jokes generated by the prompt.

### v3

# Augmenting Lateral Thinking in Language Models with Humor and Riddle Data for the BRAINTEASER Task

@@ Abstract
The SemEval 2024 BRAINTEASER task challenges language models to perform lateral thinking—a form of creative, non-linear reasoning that remains underexplored in NLP. The task comprises two subtasks, Sentence Puzzle and Word Puzzle, requiring models to defy conventional commonsense associations.
We present a system that fine-tunes DeBERTaV3 using HuggingFace’s AutoModelForMultipleChoice architecture. We augment the provided training data with two additional sources: (1) a humor-style question-answering dataset generated via GPT-4 prompting, and (2) the RiddleSense dataset. This data augmentation strategy is motivated by the observation that humor and riddles share the lateral reasoning structure required by the task. Our best system achieves 92.5% overall accuracy on the Sentence Puzzle subtask and 80.2% on the Word Puzzle subtask, ranking 6th out of 31 teams and 10th out of 23 teams, respectively. We further show that the choice of task formulation matters: framing the problem as multiple-choice rather than sequence classification yields a 10-point accuracy improvement with the same base model. Our analysis reveals that data augmentation with humor and riddle data is particularly effective for sentence-level lateral reasoning, while word-level puzzles remain a harder challenge.
Augmenting Lateral Thinking in Language Models with Humor and Riddle Data for the BRAINTEASER Task
@@ 1 Introduction
The success of language models has inspired the Natural Language Processing community to attend to tasks that require implicit and complex reasoning. Human reasoning encompasses two broad categories: lateral and vertical thinking. Lateral thinking demands unconventional, creative reasoning that deviates from traditional logical processes and has received relatively little attention from the NLP community. Vertical thinking, on the other hand, relies on logical, step-by-step reasoning and has been the focus of considerable research in recent years.
Jiang et al. (2023) introduced the BRAINTEASER dataset as a benchmark for evaluating lateral thinking in question-answering systems. The SemEval 2024 BRAINTEASER task (Jiang et al., 2024) derives a focused challenge from this dataset, comprising two subtasks: Sentence Puzzle, which requires reinterpreting sentence-level meaning, and Word Puzzle, which requires reasoning about orthographic or phonetic properties of words. In both cases, models must go beyond conventional commonsense associations and recognize non-obvious interpretations—a capability that most existing NLP benchmarks, which focus on vertical (logical, step-by-step) reasoning, do not test.
@@
Our system uses pre-trained models—BERT (Devlin et al., 2019) and DeBERTaV3 (He et al., 2023)—through HuggingFace’s (Wolf et al., 2020) AutoModelForMultipleChoice and AutoModelForSequenceClassification architectures. We mix Sentence and Word Puzzle training datasets to broaden exposure to different types of lateral thinking challenges. Additionally, we augment the training data with a humor/jokes dataset generated by GPT-4 (OpenAI et al., 2024) and the RiddleSense (Lin et al., 2021) dataset, both of which require creative, non-literal reasoning similar to the target task.
Our experiments reveal several findings. First, the AutoModelForMultipleChoice formulation substantially outperforms AutoModelForSequenceClassification, confirming that task-appropriate architecture design matters even with strong pre-trained representations. Second, augmenting training data with humor and riddle datasets consistently improves performance on the Sentence Puzzle. Third, the same augmentation strategy does not transfer to the Word Puzzle, suggesting that sentence-level and word-level lateral reasoning require different data characteristics. Our system ranked 6th out of 31 teams on the Sentence Puzzle and 10th out of 23 teams on the Word Puzzle.
Our main contributions are as follows:
  1. 1.
    We show that the multiple-choice architecture (AutoModelForMultipleChoice) significantly outperforms the sequence classification formulation for this task, yielding a 10-point accuracy gain with the same BERT base model.
  2. 2.
    We demonstrate that augmenting training data with GPT-4-generated humor QA data and the RiddleSense dataset improves lateral thinking performance, achieving 92.5% accuracy on the Sentence Puzzle.
  3. 3.
    We provide an analysis of failure modes and the differential effect of data augmentation on sentence-level versus word-level lateral reasoning.
@@
Our code and data will be available at https://github.com/soumyasmruti/semeval-2024-brainteaser after cleaning and de-anonymization.
@@ 3 Related Work
Researchers have explored various approaches to endow LLMs with commonsense reasoning abilities (Rae et al., 2021). One prominent approach is the use of knowledge graphs, which represent structured knowledge in the form of entities and their relationships (Ilievski et al., 2021). Wang et al. (2021) proposed a method for incorporating commonsense knowledge from ConceptNet (Speer et al., 2018) into language models, leading to improved performance on commonsense reasoning tasks.
Another approach involves fine-tuning pre-trained LLMs on commonsense reasoning datasets. Huang et al. (2019) introduced the COSMOS QA dataset, which consists of multiple-choice questions that require commonsense reasoning. They showed that fine-tuning pre-trained LLMs on this dataset can significantly improve their commonsense reasoning capabilities.
Researchers have also investigated the use of prompting techniques to elicit commonsense reasoning from LLMs without explicit fine-tuning. Zhou et al. (2022) proposed a method called “Conditional Prompt-Tuning” that enables LLMs to perform commonsense reasoning by conditioning on carefully designed prompts. In another work, Wei et al. (2022) showed that chain-of-thought prompting can unlock LLM reasoning ability via effective prompting techniques.
There has not been extensive research on lateral thinking in LLMs. Recently, Xie et al. (2023) proposed OlaGPT, a cognitive architecture framework that summarizes various methods of human reasoning into Chain-of-Thought (CoT) templates to improve LLM reasoning.
Data Augmentation for Reasoning.
Data augmentation has been widely used to improve model robustness in NLP tasks. In the context of reasoning, augmenting training data with related but distinct datasets has shown promise for improving generalization. Our work extends this idea by using humor and riddle datasets as augmentation sources, motivated by the structural similarity between jokes, riddles, and lateral thinking puzzles—all require reinterpreting familiar concepts in unexpected ways.
Overall, while LLMs have shown flashes of non-linear, exploratory thinking on some benchmarks, lateral thinking as a cognitive process remains an open challenge.
@@ 4.2 MultipleChoice QA with BERT and DeBERTa
4.2.1 Augmenting with RiddleSense and Humor Data
@@
The RiddleSense dataset consists of riddles—questions presented in a cryptic or metaphorical manner that challenge the reader to find a clever or unexpected answer. Riddles share a key structural property with brain teasers: they require the solver to abandon default commonsense interpretations in favor of creative alternatives.
To create the humor-style QA data, we prompted GPT-4 to generate joke-format multiple-choice questions. Jokes frequently rely on lateral thinking, as punchlines subvert the listener’s expectations by reinterpreting the setup in an unexpected way. This structural parallel to the BRAINTEASER task motivates our use of humor data as an augmentation source. The details of the data generation process are provided in Appendix A.
@@ 6.1 Subtask A : Sentence Puzzle
Initially, we trained our models only on the provided sentence puzzle dataset but soon realized that combining both the sentence puzzle and word puzzle datasets yielded better validation scores. Consequently, we used the bert-base model with AutoModelForSequenceClassification, achieving an overall accuracy of 50.8%. Given that the dataset is in a multiple-choice format, we experimented with AutoModelForMultipleChoice using the same bert model. This change significantly improved performance, increasing accuracy by 10 points to 60%. Encouraged by this, we opted for the larger DeBERTaV3 model under the AutoModelForMultipleChoice configuration. This model, combined with the original dataset, significantly boosted performance, raising overall accuracy to 88.3%. After incorporating additional datasets containing humor-style questions and the RiddleSense dataset, our best accuracy score reached 92.5%. Our approach ranked 6th among the 31 teams that participated in the task and outperformed the baseline zero shot ChatGPT by almost 50 percentage points.
@@ 6.2 Subtask B : Word Puzzle
The Word Puzzle setup followed a similar approach to the Sentence Puzzle. However, during the validation process, we found that the best-performing model was trained only on the original Word Puzzle training data combined with the humor dataset. Adding RiddleSense data and Sentence Puzzle data did not improve Word Puzzle validation scores, so we did not include those configurations in our final submission. This is a notable asymmetry: while Sentence Puzzle benefits from diverse augmentation, Word Puzzle does not. We hypothesize that this is because word-level puzzles rely on orthographic and phonetic manipulation rather than semantic reinterpretation, making sentence-oriented augmentation data less relevant. Our approach for this subtask ranked 10th among the 23 participating teams, but still outperformed the baseline zero-shot ChatGPT by approximately 40 percentage points.
6.3 Ablation Study
To isolate the contribution of each component in our system, we summarize the incremental gains on the Sentence Puzzle subtask in Table 3. Each row adds one component relative to the row above it.
The largest single gain (+28.3 points) comes from switching from BERT to DeBERTaV3, confirming the importance of the pre-trained model’s capacity. The architectural choice between sequence classification and multiple choice also yields a substantial improvement (+9.2 points), underscoring that task formulation is a meaningful design decision, not merely an implementation detail. Data augmentation provides a further 4.2-point gain on Sentence Puzzle. For Word Puzzle, the best configuration (DeBERTaV3 + AMMC + wp + Humor) achieves 80.2% overall accuracy, while adding RiddleSense or Sentence Puzzle data reduces performance.
Configuration Acc. $\Delta$
BERT + AMSC + wp + sp 50.8
BERT + AMMC wp + sp 60.0 +9.2
DeBERTaV3 + AMMC + wp + sp 88.3 +28.3
 + Humor + RS 92.5 +4.2
Table 3: Ablation on Sentence Puzzle (overall accuracy, %). RS = RiddleSense, AMSC = AutoModelForSequenceClassification, AMMC = AutoModelForMultipleChoice.
6.4 Error Analysis
We analyze the errors made by our best system to identify systematic failure patterns.
Instance-based vs. Group-based Accuracy:
A notable pattern across all configurations is the significant drop from instance-based to group-based accuracy. For example, on the Sentence Puzzle, our best system achieves 92.5% on original questions but only 87.5% on the Orig.+Sem.+Con. group metric. This indicates that while the model often answers the original question correctly, it fails on at least one adversarial variant, suggesting that its reasoning is not fully robust—it may be relying on surface patterns that break under semantic or contextual reconstruction.
Sentence Puzzle vs. Word Puzzle Gap:
Our system shows a 12.3-point gap between Sentence Puzzle (92.5%) and Word Puzzle (80.2%) overall accuracy. Sentence Puzzles require reinterpreting the semantic content of a sentence, which aligns well with the semantic representations learned by transformer models. Word Puzzles, in contrast, often require reasoning about orthographic structure (e.g., identifying that “N” is the part of “London” that is in “France”), which is not directly captured by subword tokenization. This suggests that improving Word Puzzle performance may require architectural changes or character-level augmentation strategies rather than additional semantic training data.
Effect of Augmentation on Word Puzzles:
Adding RiddleSense data, which helped the Sentence Puzzle, actually decreased Word Puzzle validation scores. RiddleSense riddles are primarily semantic in nature, which may introduce noise when the target task requires character-level or phonetic reasoning. This finding highlights that data augmentation must be carefully matched to the reasoning type required by the target task.
@@ 7 Conclusion
We presented a system for the SemEval 2024 BRAINTEASER task that achieves 92.5% accuracy on the Sentence Puzzle (6th/31 teams) and 80.2% on the Word Puzzle (10th/23 teams). Our results demonstrate three key findings: (1) the multiple-choice architecture is substantially better suited to this task than sequence classification, (2) augmenting training data with humor and riddle datasets effectively improves sentence-level lateral reasoning, and (3) data augmentation benefits do not transfer uniformly across reasoning types—word-level puzzles require different strategies. The gap between instance-based and group-based accuracy across all systems highlights that adversarial robustness remains a challenge for lateral thinking tasks. Future work should explore character-level representations or tokenization-aware approaches for word puzzles, larger-scale synthetic data generation with better quality control, and ensemble methods that combine semantic and orthographic reasoning capabilities.
8 Limitations
Our work has several limitations that should be considered when interpreting the results.
Small Training Data:
The provided training data is small (405 Sentence Puzzle and 316 Word Puzzle instances). While our augmentation strategy mitigates this to some extent, the limited data constrains the generalizability of our conclusions about which augmentation strategies are most effective.
Base-size Models Only:
We experimented exclusively with base-size models (BERT-base, DeBERTaV3-base) due to computational constraints. Larger models or recent LLM-based approaches (e.g., instruction-tuned models with in-context learning) may exhibit different augmentation dynamics and could potentially close the gap on Word Puzzles.
Synthetic Data Quality:
Our humor dataset was generated by prompting GPT-4 with relatively simple prompts and manually checked for duplicates. A more systematic generation and filtering pipeline could yield higher-quality and more diverse augmentation data. Additionally, the humor data generation process is not fully reproducible, as GPT-4 outputs are non-deterministic.
Limited Analysis of Adversarial Robustness:
While we report group-based accuracy metrics that capture robustness to semantic and context reconstruction, we do not perform a detailed analysis of which types of adversarial perturbations are most challenging. A finer-grained analysis could inform more targeted approaches to improving robustness.
Single Task Evaluation:
Our approach is evaluated solely on the BRAINTEASER task. It remains an open question whether the data augmentation strategies we propose generalize to other lateral thinking or creative reasoning benchmarks.
@@ Appendix A Humor Dataset Details
We used the following prompts to generate the humor-style dataset. We experimented with multiple prompts and gathered all outputs in a JSON file, then analyzed them manually.
@@
We initially prompted GPT-4 to generate 200 questions at once, but the output contained duplicate questions after approximately 15–16 unique ones, as the model began repeating itself. We therefore used PROMPT 3 iteratively to generate data in smaller batches, checking for duplicates manually before each addition. Below are examples of the generated jokes.